OpenStudy (anonymous):

evaluate -7 3/8-5 7/12

OpenStudy (lgbasallote):

no calculators allowed?

OpenStudy (anonymous):

\[-7 \frac{3}{8}-5\frac{7}{12}\]?????

OpenStudy (anonymous):

@shawn583

OpenStudy (anonymous):

yes I need some help

OpenStudy (anonymous):

first convert it in fractions.

OpenStudy (anonymous):

\[-(7+\frac{3}{8})-(5+\frac{7}{12})\]

OpenStudy (anonymous):

now can you solve?

OpenStudy (anonymous):

still alittle it confused on where to go from here

OpenStudy (anonymous):

take LCM

OpenStudy (anonymous):

\[-\frac{56+3}{8}-\frac{60+7}{12}\]

OpenStudy (anonymous):

so the LCD would be 24 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now what to do with the top numbers

OpenStudy (anonymous):

\[-\frac{59*3}{8*3}-\frac{67*2}{12*2}\] \[-\frac{177}{24}-\frac{134}{24}\] \[\frac{-177-134}{24}\] \[\frac{-311}{24}\]